The conjugate acid of ethanol amine (HOCH2CH2NH3+) has pKa = 9.50. What can you conclude about a solution of this amine in which [RNH3+)] = 10 x [RNH2]?)

Question 9 options:

a) pH = 10.50

b) pH = 9.50

c) pH = 8.50

d) None of the above.
Answer #1

It is basic buffer

pKb = 14-9.50 = 4.5

pOH = pKb + log([RNH3+]/[RNH2])

= 4.5+log(10/1)

= 5.5

pH = 14-5.5 = > 8.50

(c) is correct option.